# Retrieval of photon statistics from click detection

### v1

@@ Discussion.
Our comprehensive detector characterization certifies an excellent performance of the whole click detection system. The response function is mostly linear, exhibiting insignificant dark count contributions and detection nonlinearities. The overall quantum efficiency of the click-counting devices is estimated as $60\%$. Furthermore, we also ascertain the quality of the setup in terms of the uniformity of the distribution of photons across time bins without overlap, i.e., no cross-talk.
@@ IV Pseudo-photon-number reconstruction
Figure 7: Application of the Mandel $Q_{M}$ parameter [Eq. (5)] and binomial $Q_{B}$ parameter [Eq. (6)] for verifying and falsifying Poisson and binomial statistics, for $N=4$ [plots (a) and (b)] and $N=8$ detection bins [plots (c) and (d)], including a $1\sigma$ uncertainty (blue, dashed for $Q_{M}$ and purple, solid for $Q_{B}$). Plots (a) and (c) demonstrate that the measured click-counting data are compatible with a binomial distribution but distinct from a Poissonian photon-number distribution, as expected for coherent light. After pseudo-inversion, plots (b) and (d), the reconstructed statistics follow a Poisson-like, which are clearly distinct from binomial distributions, demonstrating the successful conversion from click-counting data to photon-number information.
@@
We apply the pseudo-inversion to our measured click statistics and compare the initial and resulting distribution in Fig. 6. On a superficial level, one might think that the inversion does nothing substantial. However, upon closer inspection by probing the binomial and Poissonian character in Fig. 7, we clearly find a distinction between the measured click distribution and the pseudo-photon-number distribution, which becomes increasingly prominent with increasing signal intensity of the coherent input state. In particular, we can infer that the measured data follows the expected binomial behavior [$Q_{B}=0$ for $\vec{c}\,$], and our extracted pseudo-photon values behave like a true Poisson statistics [$Q_{M}=0$ for $\vec{p}^{\,\prime}(\eta)$]. The binomial and Mandel parameter are commonly used to probe the nonclassical behavior of light. Here, we assign an additional value to such parameters: a highly sensitive and effective measure to distinguish quantum-statistical signatures.
@@ VI Conclusion
In this work, a pseudo-inversion of click-counting data was carried out to obtain the otherwise inaccessible photon-number distribution. The Mandel and binomial parameters were proposed and used as measures to assess the quality of the reconstruction procedure, and to generally discern different types of quantum statistics.
@@
Our experimental implementation of the methodology is based on a custom-made, time-bin multiplexing detector that is easily configurable to a desired resolution. Based on detector tomography with coherent light of different intensities, we analyzed the detection system’s response, allowing us to determine essential properties, such as detection efficiency, cross-talk between time bins, etc. To demonstrate the variable nature of our setup, both a four-bin and eight-bin configuration were analyzed in this manner. Using the binomial and Mandel parameter, we proved that our coherent-light data are consistent with a binomial click distribution and inconsistent with a Poissonian photon distribution as one would expect from the comparison of click and photoelectric counting theories.
@@
A pseudo-inversion approach was then implemented to recover a pseudo-photon-number resolution up to four and eight photons, depending on the detector’s configuration. While the resulting and initially measured distribution appear rather similar when plotted together, the Mandel and binomial parameters reveal the now Poissonian nature of the reconstructed distribution and rule out the previous binomial behavior, respectively. Thereby, the success of the pseudo-inversion as well as the effectiveness of the discrimination via the two parameters is demonstrated. Importantly, the pseudo-inversion is based on an analytically obtained matrix that is applied only once to each data set, circumventing difficulties and computational complexities of other competing reconstruction techniques.
@@
Eventually, a loss deconvolution was applied to correct for the previously estimated losses of the detection system. The resulting distribution is virtually indistinguishable from the Poisson distribution of a perfect (e.g., lossless) photon detector, up to the threshold photon number given by the maximal number of total clicks (here, four and eight). To enable a fair comparison between reconstructed and ideal statistics, the mean photon number of the coherent input states were measured independently from the click-counting device.
@@
For determining nonclassical properties of quantum light, a thorough detector calibration and readily accessible reconstruction of quantum features, as laid out in this work, is essential. Thereby, experimentally accessible click-counting devices become one of the closest, readily available substitute to true photon-number-resolving detectors. Furthermore, an extra application of parameters, which are commonly used to determine nonclassicality in quantum optics, was provided by exploiting them to discern fundamentally different quantum statistics in a highly sensitive manner. Also, the application of the binomial and Mandel parameters shows that a broad belief—a click corresponds to a photon—cannot be true because of their distinct binomial and Poissonian behavior, respectively. In the future, we plan to extend our framework to emerging threshold detectors, offering an intrinsic pseudo-photon-number resolution, e.g., in Refs. [79, 80], as well as generalization to phase-sensitive detector applications [81, 82].

### v2

@@ Discussion.
Our comprehensive detector characterization certifies an excellent performance of the whole click detection system. The response function is mostly linear, exhibiting insignificant dark count contributions and detection nonlinearities. The overall quantum efficiency of the click-counting devices is estimated as $60\%$. Furthermore, we also ascertain the quality of the setup in terms of the uniformity of the distribution of photons across time bins without overlap, i.e., no cross talk.
@@ IV Pseudo-photon-number reconstruction
Figure 7: Application of the Mandel $Q_{M}$ parameter [Eq. (2)] and binomial $Q_{B}$ parameter [Eq. (3)] for verifying and falsifying Poisson and binomial statistics, for $N=4$ [plots (a) and (b)] and $N=8$ detection bins [plots (c) and (d)], including a $1\sigma$ uncertainty (blue, dashed for $Q_{M}$ and purple, solid for $Q_{B}$). Plots (a) and (c) demonstrate that the measured click-counting data are compatible with a binomial distribution but distinct from a Poissonian photon-number distribution, as expected for coherent light. After pseudo-inversion, plots (b) and (d), the reconstructed statistics follow Poisson-like distributions, which are clearly distinct from binomial distributions, demonstrating the successful conversion from click-counting data to photon-number information.
@@
We apply the pseudo-inversion to our measured click statistics and compare the initial and resulting distribution in Fig. 6. On a superficial level, one might think that the inversion does nothing substantial. However, upon closer inspection by probing the binomial and Poissonian character in Fig. 7, we clearly find a distinction between the measured click distribution and the pseudo-photon-number distribution, which becomes increasingly prominent with increasing signal intensity of the coherent input state. In particular, we can infer that the measured data follow the expected binomial behavior [$Q_{B}=0$ for $\vec{c}\,$], and our extracted pseudo-photon values behave like a true Poisson statistics [$Q_{M}=0$ for $\vec{p}^{\,\prime}(\eta)$]. The binomial and Mandel parameter are commonly used to probe the nonclassical behavior of light. Here, we assign an additional value to such parameters: a highly sensitive and effective measure to distinguish quantum-statistical signatures.
@@ VI Conclusion
In this paper, a pseudo-inversion of click-counting data was carried out to obtain the otherwise inaccessible photon-number distribution. The Mandel and binomial parameters were proposed and used as measures to assess the quality of the reconstruction procedure, and to generally discern different types of quantum statistics.
@@
Our experimental implementation of the methodology is based on a custom-made, time-bin multiplexing detector that is easily configurable to a desired resolution. Based on detector tomography with coherent light, we analyzed the detection system’s response, allowing us to determine essential properties, such as detection efficiency and cross-talk between time bins. To demonstrate the variable nature of our setup, both a four-bin and eight-bin configuration were analyzed in this manner. Using the binomial and Mandel parameter, we proved that our coherent-light data are consistent with a binomial click distribution and inconsistent with a Poissonian photon distribution as one would expect from the comparison of click and photoelectric counting theories.
@@
A pseudo-inversion approach was implemented to recover a pseudo-photon-number resolution up to four and eight photons, depending on the detector’s configuration. While the resulting and initially measured distribution appear rather similar when plotted together, the Mandel and binomial parameters reveal the now Poissonian nature of the reconstructed distribution and rule out the previous binomial behavior, respectively. Thereby, the success of the pseudo-inversion as well as the effectiveness of the discrimination via the two parameters is demonstrated. Importantly, the pseudo-inversion is based on an analytically obtained matrix that is applied only once to each data set, circumventing difficulties and computational complexities of other competing reconstruction techniques. The reconstruction method was applied to coherent states at the boundary between classical and nonclassical light, presenting a uniquely challenging scenario. However, it is worth mentioning that the pseudo-inversion applies to general, unknown states.
@@
Eventually, a loss deconvolution was applied to correct for the previously estimated losses of the detection system. The resulting distribution is virtually indistinguishable from the Poisson distribution of a perfect (e.g., lossless) photon detector, up to the threshold photon number given by the maximal number of total clicks (here, four and eight). To enable a fair comparison between reconstructed and ideal statistics, the mean photon number of the coherent input states was measured independently from the click-counting device.
@@
For determining nonclassical properties of quantum light, a thorough detector calibration and readily accessible reconstruction of quantum features, as laid out in this work, is essential. Thereby, experimentally accessible click-counting devices become one of the closest, readily available substitutes to true photon-number-resolving detectors. Furthermore, an extra application of parameters, which are commonly used to determine nonclassicality in quantum optics, was provided by exploiting them to discern fundamentally different quantum statistics in a highly sensitive manner. Also, the application of the binomial and Mandel parameters shows that a broad belief—a click corresponds to a photon—cannot be true because of their distinct binomial and Poissonian behavior, respectively. In the future, we plan to extend our framework to emerging threshold detectors, offering an intrinsic pseudo-photon-number resolution, e.g., in Refs. [60, 61], as well as generalization to phase-sensitive detector applications [62, 63].